Find the vertex and intercept only.
y=3-5x+2x^2
I would assume to put them in order, but I'm not really sure from there. Help is Appreciated thanks! :)

Answer by stanbon(75887)

Find the vertex and intercept only.
y=3-5x+2x^2
-----
y = 2x^2 - 5x + 3
-----
2x^2-5x = y-3
----
2(x^2-(5/2) + (5/4)^2) = y - 3 + 2(5/4)^2
----
2(x+(5/4))^2 = y + 9/64
------
Vertex: (-5/4, -9/64)
---
y-intercept:
Let x = 0, then y = 3
=========================
Cheers,
Stan H.
=========================